OpenStudy (anonymous):

On a particular day, the wind added 3 km per hour to Alfonso’s rate when he was cycling with the wind and subtracted 3 km per hour from his rate on his return trip. Alfonso found that in the same amount of time he could cycle 36.5 km with the wind, he could go only 24 km against the wind. What is his normal bicycling speed with no wind? a. 15 km/h c. 35 km/h b. 40 km/h d. 45 km/h

OpenStudy (anonymous):

ldk c

OpenStudy (anonymous):

let his normal bicycling speed be v km/h then with the wind it becomes v+3 and against the wind it becomes v-3 as we know distance traversed varies directly with speed if time is constant then we have as per the question (v+3)/(v-3) =36.5/24 now we need to find v

OpenStudy (anonymous):

now applying componendo and dividendo we have (v+3+v-3)/(v+3-v+3) =(36.5 +24)/(36.5-24) or v/3 =60.5/12.5 or v = 14.52 km/h (appx) so may be a. 15 km/h